OpenStudy (anonymous):

Suppose a cube of aluminum which is 1.00 cm on a side accumulates a net charge of +3.00 pC. What is the percentage of the electrons originally in the cube was removed? Here is a list of fun facts I collected: Atomic weight of Aluminum 26.98 g/ mol Atoms per mole \[6.022\times 10^{23}\] Density \[2.70g\cdot cm^3\] Electric charge \[-1.60\times 10^{-19}\;C\] Mass of an electron \[9.11\times 10^{-31}\;kg\]

OpenStudy (anonymous):

\[1\; cm^3 \cdot \frac{2.70g}{cm^3}\cdot \frac{mol}{26.98\;g}\cdot \frac{6.022\times10^{23}\;atoms}{mol}\cdot \frac{13\; electrons}{atom}\cdot \frac{1.60\times10^{-19}C}{electron}\]

OpenStudy (anonymous):

\[\small{1\; \cancel{cm^3} \cdot \frac{2.70\cancel{g}}{\cancel{cm^3}}\cdot \frac{\cancel{mol}}{26.98\;\cancel{g}}\cdot \frac{6.022\times10^{23}\;\cancel{atoms}}{\cancel{mol}}\cdot \frac{13\; \cancel{electrons}}{\cancel{atom}}\cdot \frac{1.60\times10^{-19}C}{\cancel{electron}}}\]

OpenStudy (anonymous):

125350.4522 C

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

\[3pC\cdot \frac{C}{1\times10^{-12}\; pC}\cdot \frac{electron}{1.60\times10^{-19}\;C}\]

OpenStudy (anonymous):

\[3pC\cdot \frac{C}{1\times10^{-12}\; pC}\cdot \frac{electron}{1.60\times10^{-19}\;C}=1.875\times10^{31}\; electrons\] This must be the number of electrons that were removed...

OpenStudy (anonymous):

(Electrons originally-electrons removed)/electrons originally *100 \[\frac{(7.8344E23)-(1.875E31)}{7.8344E23}*100=1.87E33\] What am I doing wrong?

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

the example attached used 2.5pC instead of 3...but it's the same idea

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

I got it. \[\frac{accumulated\;charge\cdot Atomic\;weight}{charge\;e \cdot\left(\frac{g\; Al \;in \;1\;cm^3}{mol\; Al \; in\;1\;cm^3}\right)}\cdot 100= \]